OpenStudy (anonymous):

Can someone help me with this please? I don't know how to do this. Write the equation for the horizontal line that contains point G(-8, 8).

OpenStudy (anonymous):

Horizontal lines are of the form y=n where you may have been given a point in the form (m, n). You have such a point G.

OpenStudy (anonymous):

keep in mind that a horizontal line is parallel to x axis, and slope is the value of tan of the angle between the x axis and the line

OpenStudy (anonymous):

@piscez.in: What do you mean by the slope is the value of tan of the angle between the x-axis and the line? Specifically, what do you mean by the value of tan? @tcarroll010: Okay, just a moment please.

OpenStudy (anonymous):

So it would be something like y = 8 - (-8)?

OpenStudy (anonymous):

@fang02 tan value as in tangent value, like sine cosine, tan etc.

OpenStudy (anonymous):

Hmm..I do not believe I have learned about the tan value then.

OpenStudy (anonymous):

He means that the equation of a line in slope-intercept form, in general is y = mx + b and that here we have "0" slope, so the mx term disappears. So, we have y=b, or in my notation, y=n. You have point (-8, 8) which is the same distance away from the x-axis as the "b" intercept, (0, 8), so you can simplify and use the "n".

OpenStudy (anonymous):

@Fang02 oh im sorry hun, well then forget it. But try to roughly draw the line on a paper. Find the y intercept. And you must already be knowing that the slope of horizontal lines( lines parallel to x axis) is always 1

OpenStudy (anonymous):

Not y = 8 - (-8). Instead, y=n where n = 8

OpenStudy (anonymous):

So, the answer is pretty much y = 8 then? No need to be sorry pisez.in.

OpenStudy (anonymous):

Yes! You got it. Good job! It's just y=8.

OpenStudy (anonymous):

Okay, thanks a lot! :) Would you mind helping me with another problem real quick?

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(4, 1) with slope -4.

OpenStudy (anonymous):

Point-slope form is y2 - y1 over x2 over x1 So, I think it would be: 1 - 4 over something

OpenStudy (anonymous):

point-slope form is going to be:\[y - y _{1} = m(x - x _{1})\] Here, y is y and x is x as variables. x1 and y1 are a SPECIFIC point in the form (x1, y1)

OpenStudy (anonymous):

(x1, y1) is the point you are given that is (4, 1), so x1 is 4 and y1 is 1.

OpenStudy (anonymous):

I think that (x1, y1) would be (-4, 1) right?

OpenStudy (anonymous):

Why does the 4 turn positive?

OpenStudy (anonymous):

(x1, y1) is the point you are given that is (4, 1), so x1 is 4 and y1 is 1. It's the slope that is -4.

OpenStudy (anonymous):

I understand.

OpenStudy (anonymous):

So now, just substitute.

OpenStudy (anonymous):

y - y1 = m(x - x1) y - 1 = -4 (4 - 4) ?

OpenStudy (anonymous):

Not quite there yet. What happened to "x"? "x" stays as "x" just like "y" stays as "y". You only substitute for y1, x1, and m.

OpenStudy (anonymous):

Oh okay. So, y = 1 = -4(x - 4)

OpenStudy (anonymous):

You still have a small typo, but you are extremely close now.

OpenStudy (anonymous):

Would I put x1 in parenthesis to show that it is positive and is just being subtracted? y = 1 = -4(x - (4))

OpenStudy (anonymous):

You don't have to do that, you can just leave it as y - 1 = -4(x - 4). Notice how after the "y" I changed that from "=" to "-"

OpenStudy (anonymous):

Oh okay, I understand now.

OpenStudy (anonymous):

Very nice working with you and you will get there. Just keep up the good work!

OpenStudy (anonymous):

Thanks! You too! :)